# Discovering overlapping communities in multi-layer directed networks

### v1

@@ Abstract
This article explores the challenging problem of detecting overlapping communities in multi-layer directed networks. Our goal is to understand the underlying asymmetric overlapping community structure by analyzing the mixed memberships of nodes. We introduce a new model, the multi-layer mixed membership stochastic co-block model (multi-layer MM-ScBM), to model multi-layer directed networks in which nodes can belong to multiple communities. We develop a spectral procedure to estimate nodes’ memberships in both sending and receiving patterns. Our method uses a successive projection algorithm on a few leading eigenvectors of two debiased aggregation matrices. To our knowledge, this is the first work to detect asymmetric overlapping communities in multi-layer directed networks. We demonstrate the consistent estimation properties of our method by providing per-node error rates under the multi-layer MM-ScBM framework. Our theoretical analysis reveals that increasing the overall sparsity, the number of nodes, or the number of layers can improve the accuracy of overlapping community detection. Extensive numerical experiments are conducted to validate these theoretical findings. We also apply our method to one real-world multi-layer directed network, gaining insightful results.
@@ 1 Introduction
Community detection is a powerful tool for learning the latent community structure within networks. Its core objective is to extract community information of nodes solely based on the network’s topology. For a comprehensive understanding of this technique, refer to [2, 3], which provide detailed introductions to the field of community detection. Notably, in real-world networks, a node can often be affiliated with multiple communities. In the context of directed networks, where interactions between nodes are asymmetric and directional, each node exhibits two distinct community patterns: a sending pattern and a receiving pattern. The overlapping nature of nodes, their asymmetric relationships, and the multi-layer characteristics of multi-layer directed networks further complicate community detection. We focus on the challenging problem of discovering asymmetric overlapping communities in multi-layer directed networks in this article.
@@
A widely studied variant of this problem focuses on the scenario where the network is undirected, and each node exclusively belongs to a single community. For example, the multi-layer stochastic block model (multi-layer SBM), as explored in various studies [4, 5, 6, 7], generates each layer of a multi-layer undirected network from the stochastic block model (SBM) [8], a popular statistical model for single-layer undirected networks in which communities are mutually exclusive. [4] demonstrated the estimation consistency of a spectral approach designed based on the sum of adjacency matrices, specifically under conditions where the number of layers increases while the number of nodes remains fixed, within the framework of the multi-layer SBM. [5] conducted a study on the theoretical guarantees of spectral and matrix factorization methods within the context of the multi-layer SBM. [6] introduced a least squares estimation method for communities and demonstrated its estimation consistency within the multi-layer SBM. Notably, the debiased spectral clustering algorithm introduced in [7], which offers theoretical guarantees, significantly outperforms the spectral method using the sum of adjacency matrices, the Linked Matrix Factorization method, and the Co-regularized Spectral Clustering method studied in [5]. Also see [9, 10, 11, 12] for other recent works in multi-layer undirected networks in which communities are disjoint. Unluckily, applying all algorithms above to discover communities in multi-layer directed networks is not feasible. To address this limitation, [1] proposed a spectral co-clustering algorithm that leverages the debiased approach introduced in [7] to identify disjoint communities under the multi-layer stochastic co-block model (multi-layer ScBM). This model extends the stochastic co-block model (ScBM) presented in [13] from single-layer directed networks to multi-layer directed networks. While the works in [1] address the asymmetric and directional characteristics in multi-layer directed networks, they do not extend to the general problem addressed in this article, where a node may simultaneously belong to more than one community. The key contributions of this article are as follows:
@@
Notation: For any positive integer $q$, let $[q]$ be the set $\{1,2,\ldots,q\}$, $I_{q\times q}$ be the $q\times q$ identity matrix, and $\|x\|_{q}$ be the $l_{q}$ norm of vector $x$. For any matrix $M$, we use $M^{\prime}$ to denote its transpose, $\kappa(M)$ for its condition number, $\|M\|_{F}$ for the Frobenius norm, $\mathrm{rank}(M)$ for its rank, $M(S,:)$ for the submatrix comprising the rows of $M$ indexed by set $S$, $\lambda_{k}(M)$ for the $k$-th largest eigenvalue (ordered in absolute value), $\|M\|_{\infty}$ for the maximum absolute row sum, and $\|M\|_{2\rightarrow\infty}$ for the maximum row-wise $l_{2}$ norm. We further denote expectation with $\mathbb{E}[\cdot]$ and probability with $\mathbb{P}(\cdot)$. Additionally, the phrase “leading $K$ eigenvectors” refers to the eigenvectors respective to the top $K$ eigenvalues ordered in absolute value. Lastly, $e_{i}$ represents the standard basis vector that has the $i$-th element equal to 1, while all other elements are set to zero.
2 Related works
@@
Our proposed method has a close relationship with the debiased spectral clustering methodologies presented in [7, 1]. For non-overlapping community detection in multi-layer undirected networks, [7] proposes a debiased spectral clustering algorithm by applying the K-means algorithm to several leading eigenvectors of an aggregation matrix under the multi-layer SBM model. [1] extends the idea in [7] from multi-layer undirected networks modeled by multi-layer SBM to multi-layer directed networks modeled by multi-layer ScBM. The numerical findings in [7] and [1] indicate that debiased spectral clustering achieves superior performance compared to traditional spectral clustering methods that rely on the summation of adjacency matrices or the squared summation of adjacency matrices.
@@
However, multi-layer ScBM can not capture the overlapping community structure of multi-layer directed networks. The mixed membership stochastic block model (MMSB) [15] allows each node to belong to multiple communities for single-layer undirected networks. The multi-way blockmodel introduced in [16] extends MMSB from undirected networks to directed networks. [17] develops an efficient spectral algorithm to estimate overlapping community memberships and derives the node-wise error bounds under MMSB. [18] develops an algorithm with theoretical guarantees to fit a degree-corrected version of MMSB. [19] designs two regularized spectral clustering algorithms based on a regularized Laplacian matrix to estimate nodes’ memberships under MMSB. For works on overlapping community detection in [17, 19, 18], vertex hunting algorithms developed in [20, 14, 21, 18] are applied to hunt for pure nodes, as defined later. In this article, the proposed spectral method combines the debiased idea in [7, 1] and the vertex hunting technique in [17] to estimate nodes’s mixed memberships under the proposed model.
@@
For our theoretical analysis, we obtain the row-wise eigenspace error of two debiased aggregation matrices by using the Bernstein inequality of [22] and Theorem 4.2 of [23]. Subsequently, we employ theoretical frameworks from [17] to establish row-wise error bounds. By doing so, we broaden the scope of the theoretical analysis in [17], extending it from single-layer undirected networks to multi-layer directed networks, and from adjacency matrices to debiased aggregation matrices.
@@ 3 Model framework: multi-layer MM-ScBM
[…]
(a) Layer 1
[…]
(b) Layer 2
[…]
(c) Layer 3
[…]
(d) Layer 4
[…]
(e) Layer 5
[…]
(f) Layer 6
@@
This article assumes that the multi-layer directed network $\mathcal{N}$ comprises $L$ distinct layers, each sharing a common set of $n$ interconnected nodes. Let $A_{l}\in\{0,1\}^{n\times n}$ denote the adjacency matrix of the $l$-th layer directed network $\mathcal{N}_{l}$, where $A_{l}(i,j)=1$ indicates the presence of a directed edge from node $i$ to node $j$ (i.e., $i\rightarrow j$) and $A_{l}(i,j)=0$ otherwise, for all $i,j\in[n]$ and $l\in[L]$. Therefore, in $A_{l}$, its $i$-th row captures the outgoing edges from node $i$, while its $i$-th column details the incoming edges to node $i$ at the $l$-th layer for $i\in[n],l\in[L]$. Figure 1 displays a toy example of a multi-layer directed network, showcasing how nodes establish distinct connections across various layers. For the sending pattern, we assume that all nodes are classified into $K$ distinguishable row communities:
$\mathcal{C}_{r,1},\mathcal{C}_{r,2},\ldots,\mathcal{C}_{r,K}.$ (1)
Let $\Pi_{r}$ be an $n\times K$ row membership matrix that is common to all layers. Specifically, $\Pi_{r}(i,k)$ represents the “weight” indicating the affiliation of node $i$ to the $k$-th row community $\mathcal{C}_{r,k}$ for $i\in[n],k\in[K]$.
@@
$\mathcal{C}_{c,1},\mathcal{C}_{c,2},\ldots,\mathcal{C}_{c,K}.$ (2)
Furthermore, let $\Pi_{c}$ be an $n\times K$ column membership matrix that is common across all layers and $\Pi_{c}(i,k)$ represents the “weight” indicating the affiliation of node $i$ to the $k$-th column community $\mathcal{C}_{c,k}$ for $i\in[n],k\in[K]$.
@@
each row and column community must contain at least one pure node. (3)
Without loss of generality, we rearrange the nodes’ order such that $\Pi_{r}(\mathcal{I}_{r},:)=I_{K\times K}$ and $\Pi_{c}(\mathcal{I}_{c},:)=I_{K\times K}$. Here, $\mathcal{I}_{r}$ represents the indices of $K$ pure nodes, each belonging to a distinct row community, while $\mathcal{I}_{c}$ denotes the indices of $K$ pure nodes, each from a unique column community. For each adjacency matrix $A_{l}$ for $l\in[L]$, the model considered in this article assumes that it is generated from a directed version of MMSB. In particular, we consider the following model.
@@
where $B_{l}\in[0,1]^{K\times K}$ and $\rho\in(0,1]$.
In the definition of multi-layer MM-ScBM, the asymmetric matrix $B_{l}$ can be different for different layers, and this guarantees that $A_{l}$ may be generated from the model with common membership matrices and possibly different edge probabilities. Meanwhile, since $\mathbb{P}(A_{l}(i,j)=1)=\Omega_{l}(i,j)=\rho\Pi_{r}(i,:)B_{l}\Pi^{\prime}_{c}(j,:)\leq\rho$ for $i\in[n],j\in[n],l\in[L]$, we observe that a reduction in the value of $\rho$ leads to a decrease in the likelihood of an edge forming between any two nodes, thereby controlling the overall sparsity of the multi-layer bipartite network. Consequently, we refer to $\rho$ as the sparsity parameter in this article. Given the set of $L$ adjacency matrices $\{A_{l}\}^{L}_{l=1}$, our goal is to estimate $\Pi_{r}$ and $\Pi_{c}$ in this article. Similar to the consistency analysis conducted in [24, 17] for single-layer networks, we will permit $\rho$ to approach zero as $n$ and $L$ increase in our theoretical analysis.
@@
Our model encompasses numerous previous models as specific cases. When all nodes are pure, it simplifies to the multi-layer ScBM discussed in [1]. If the network is undirected and all nodes are pure, it reduces to the multi-layer SBM studied in [4, 5, 6, 7]. Further, for $L=1$, it degenerates to the multi-way blockmodel presented in [16]. In the case of $L=1$ and an undirected network, it reduces to the MMSB model proposed in [15]. Moreover, when all nodes are pure and the network only has a single layer, it reduces to the ScBM model described in [13]. Finally, if the network is also undirected, it simplifies to the SBM introduced in [8].

### v2

@@ Abstract
Community detection in multi-layer undirected networks has attracted considerable attention in recent years. However, multi-layer directed networks are common in the real world, and existing community detection methods often either ignore the asymmetric structure in multi-layer directed networks or assume that every node solely belongs to a single community, significantly limiting their applicability to overlapping multi-layer directed networks, where nodes can belong to multiple communities simultaneously. To fill this gap, this article explores the challenging problem of detecting overlapping communities in multi-layer directed networks. Our goal is to understand the underlying asymmetric overlapping community structure by analyzing the mixed memberships of nodes. We introduce a novel multi-layer mixed membership stochastic co-block model (multi-layer MM-ScBM) to model overlapping multi-layer directed networks. We develop a spectral procedure to estimate nodes’ memberships in both sending and receiving patterns. Our method uses a successive projection algorithm on a few leading eigenvectors of two debiased aggregation matrices. To our knowledge, this is the first work to detect asymmetric overlapping communities in multi-layer directed networks. We demonstrate the consistent estimation properties of our method by providing per-node error rates under the multi-layer MM-ScBM framework. Our theoretical analysis reveals that increasing the overall sparsity, the number of nodes, or the number of layers can improve the accuracy of overlapping community detection. Extensive numerical experiments validate these theoretical findings. We also apply our method to one real-world multi-layer directed network, gaining insightful results.
@@ 1 Introduction
Community detection is a powerful tool for learning the latent community structure within networks. Its core objective is to extract community information of nodes solely based on the network’s topology. For a comprehensive understanding of this technique, refer to [2, 3, 4], which provide detailed introductions to the field of community detection. Notably, in real-world networks, a node can often be affiliated with multiple communities [5, 6, 7, 8]. In the context of directed networks, where interactions between nodes are asymmetric and directional, each node exhibits two distinct community patterns: a sending pattern and a receiving pattern [9, 1]. The overlapping nature of nodes, their asymmetric relationships, and the multi-layer characteristics of multi-layer directed networks further complicate community detection. We focus on the challenging problem of discovering asymmetric overlapping communities in multi-layer directed networks in this article.
Table 1: A brief summary of community detection methods in multi-layer networks. Abbreviations: MLE (Maximum Likelihood Estimation), LSE (Least Squares Estimation), and NMF (Nonnegative Matrix Factorization).
Methods considered Methods type Directed networks? Overlapping communities? Theoretical guarantees?
Spectral Mean and Profile MLE [10] Spectral-based and MLE-based No No Yes
S2-jNMF [11] NMF-based No No No
CSNMF and CSNMTF [12] NMF-based No No No
SCDSBM [13] Spectral-based No No Yes
IterModMax [14] Modularity-based No No No
MLCD [15] Multi-objective-optimization-based No Yes No
Multimodbp [16] Modularity-based No No No
LSE [17] LSE-based No No Yes
OLMF, Mean adj, Coreg spec, and SpecK [18] Matrix-factorization-based and spectral-based No No Yes
Cr-ENMF [19] NMF-based No No No
UASE [20] Spectral-based No No Yes
TWIST [21] Tensor-based No No Yes
MASE [22] Spectral-based No No Yes
MjNMF [23] NMF-based No No No
NE2NMF [24] NMF-based No No No
MA, MVM, and MVP [25] Modularity-based No No No
ALMA [26] Tensor-based No No Yes
SSJSNMF [27] NMF-based No No No
SoS-Debias [28] Spectral-based No No Yes
CAMSBM [29] Tensor-based No No Yes
ISC [30] Spectral-based No No Yes
HDP-LPCM [31] Bayesian-based No No No
McNMF [32] NMF-based No No No
CDMA [33] Motif-based No No No
MOEA-CPI [34] Multi-objective-evolutionary-based No No No
CDMMHN [35] Motif-modularity-based No No No
WPTDD [36] Tensor-based No No No
HSBM [37] Bayesian-based No No No
DIMPLE-GDPG [38] Tensor-based No No Yes
MX-ONMTF [39] NMF-based No No No
SPATMME [40] Tensor-based No Yes Yes
MCLCD [41] NMF-based No No No
NMFTD [42] NMF&\&&Tensor-based No Yes Yes
NSoA and NDSoSA [43] Spectral-based No No Yes
DSoG [1] Spectral-based Yes No Yes
CSPDSoS (ours) Spectral-based Yes Yes Yes
@@
A widely studied variant of this problem focuses on the scenario where the network is undirected, and each node exclusively belongs to a single community. For example, the multi-layer stochastic block model (multi-layer SBM), as explored in various studies [10, 18, 17, 28], generates each layer of a multi-layer undirected network from the stochastic block model (SBM) [44], a popular statistical model for single-layer undirected networks in which communities are mutually exclusive. [10] demonstrated the estimation consistency of a spectral approach designed based on the sum of adjacency matrices, specifically under conditions where the number of layers increases while the number of nodes remains fixed, within the framework of the multi-layer SBM. [18] conducted a study on the theoretical guarantees of spectral and matrix factorization methods within the context of the multi-layer SBM. [17] introduced a least squares estimation method for communities and demonstrated its estimation consistency within the multi-layer SBM. Notably, the debiased spectral clustering algorithm introduced in [28], which offers theoretical guarantees, significantly outperforms the spectral method using the sum of adjacency matrices, the Linked Matrix Factorization method, and the Co-regularized Spectral Clustering method studied in [18]. Also see [21, 22, 26, 29] for other recent works in multi-layer undirected networks in which communities are disjoint.
Unluckily, applying all aforementioned algorithms to discover communities in multi-layer directed networks is not feasible. To address this limitation, [1] proposed a spectral co-clustering algorithm that leverages the debiased approach introduced in [28] to identify disjoint communities under the multi-layer stochastic co-block model (multi-layer ScBM). This model extends the stochastic co-block model (ScBM) presented in [9] from single-layer directed networks to multi-layer directed networks. While the works in [1] address the asymmetric and directional characteristics in multi-layer directed networks, they do not extend to the general problem addressed in this article, where a node may simultaneously belong to more than one community. Recently, [40, 42] introduced tensor-based methods with theoretical guarantees for estimating mixed memberships in overlapping multi-layer undirected networks. Furthermore, various other approaches for community detection in multi-layer networks exist, such as NMF-based and modularity-based techniques, as summarized in Table 1. Spectral-based methods, such as those proposed in [10, 28, 1], which rely on the eigendecomposition or singular value decomposition of certain matrices, are also included in this table. However, as Table 1 clearly indicates, all previous methods either focus on non-overlapping communities or are tailored for multi-layer undirected networks. To the best of our knowledge, there is currently no efficient method available for estimating nodes’ mixed memberships in overlapping multi-layer directed networks, where the relationships between nodes are directional and each node can belong to multiple communities with varying weights. This article aims to fill this gap. The key contributions of this article are as follows:
@@
Notation: For any positive integer $q$, let $[q]$ be the set $\{1,2,\ldots,q\}$, $I_{q\times q}$ be the $q\times q$ identity matrix, and $\|x\|_{q}$ be the $l_{q}$ norm of vector $x$. For any matrix $M$, we use $M^{\prime}$ to denote its transpose, $\kappa(M)$ for its condition number, $\|M\|_{F}$ for the Frobenius norm, $\mathrm{rank}(M)$ for its rank, $M(S,:)$ (and $M(:,S)$) for the submatrix comprising the rows (and columns) of $M$ indexed by set $S$, $\lambda_{k}(M)$ for the $k$-th largest eigenvalue (ordered in absolute value), $\|M\|_{\infty}$ for the maximum absolute row sum, and $\|M\|_{2\rightarrow\infty}$ for the maximum row-wise $l_{2}$ norm. We further denote expectation with $\mathbb{E}[\cdot]$ and probability with $\mathbb{P}(\cdot)$. Additionally, the phrase “leading $K$ eigenvectors” refers to the eigenvectors respective to the top $K$ eigenvalues ordered in absolute value. Lastly, $e_{i}$ represents the standard basis vector that has the $i$-th element equal to 1, while all other elements are set to zero.
2 Related work
The field of community detection in multi-layer networks has garnered significant attention, with various approaches being proposed in the literature to address the complexity of identifying latent community structures within such networks. In this section, we provide a structured discussion of relevant works, highlighting their connections and differences with respect to our work in this article.
2.1 Spectral-based methods
Recall that Table 1 summarizes substantial existing works, including spectral-based, NMF-based, and modularity-based methods, for community detection in multi-layer networks. However, only a few works listed in this table are closely related to this study because this article focuses specifically on spectral-based methods due to their well-known efficiency, ease of implementation, and feasibility for theoretical analysis [46, 47, 48].
@@ 2.1 Spectral-based methods
Within the spectral-based methods for multi-layer networks, our proposed method has a close relationship with the debiased spectral clustering methodologies presented in [28, 1, 43]. For non-overlapping community detection in multi-layer undirected networks, [28] initially proposes a debiased spectral clustering algorithm by applying the K-means algorithm to several leading eigenvectors of an aggregation matrix under the multi-layer SBM model. [1] extends the idea in [28] from multi-layer undirected networks modeled by multi-layer SBM to multi-layer directed networks modeled by multi-layer ScBM. [43] extends works in [28] by considering nodes’ degree variations in the real world, enhancing the debiased spectral clustering framework. The numerical findings in [28, 1, 43] indicate that debiased spectral clustering achieves superior performance compared to traditional spectral clustering methods that rely on the summation of adjacency matrices or the squared summation of adjacency matrices. In particular, [43] theoretically proved the superiority of the debiased spectral method over the spectral mean method [10]. However, works in [28, 1, 43] are limited to non-overlapping communities, failing to capture the overlapping community structure prevalent in real-world networks.
2.2 Overlapping community detection
While the spectral-based methods proposed in [28, 1, 43] have been successful in detecting non-overlapping communities in multi-layer networks, the task of identifying overlapping communities is more challenging. The mixed membership stochastic block model (MMSB) [5] allows each node to belong to multiple communities for single-layer undirected networks, and it is one of the most popular statistical models in network science due to its generality, flexibility, and interpretability. The multi-way blockmodel introduced in [49] extends MMSB from undirected networks to directed networks.
@@ 2.2 Overlapping community detection
Spectral-based methods have also been developed for overlapping community detection in single-layer networks. For example, based on the observation that there exists an ideal simplex structure inherent in the leading eigenvector matrix of the expectation adjacency matrix, [8] develops an efficient spectral algorithm to estimate overlapping community memberships and derives the node-wise error bounds under MMSB. [50] develops an algorithm called Mixed-SCORE with theoretical guarantees to fit a degree-corrected version of MMSB, where Mixed-SCORE is also designed based on an ideal simplex structure inherent in the row-wise ratio matrix of the leading eigenvector matrix from the expectation adjacency matrix. The Mixed-SCORE algorithm can be viewed as an extension of the SCORE algorithm developed in [51] from non-overlapping networks to overlapping networks. [52] designs two regularized spectral clustering algorithms based on a regularized Laplacian matrix to estimate nodes’ memberships under MMSB. For works on overlapping community detection in [8, 52, 50], vertex hunting algorithms developed in [53, 45, 54, 50] are applied to hunt for pure nodes, as defined later.
2.3 Connections to previous work
Despite the progress made in overlapping community detection, existing methods are primarily designed for single-layer networks. To address the limitations of existing approaches, we first propose a novel statistical model multi-layer MM-ScBM to model overlapping communities in multi-layer directed networks. This model extends the MMSB and multi-way blockmodel from single-layer networks to multi-layer directed networks, capturing the asymmetric and directional relationships between nodes. We then develop a spectral-based method to estimate nodes’ mixed memberships in both the sending and receiving patterns under the multi-layer MM-ScBM framework. Our method integrates the debiased idea in [28, 1, 43] with the vertex hunting technique in [8] to estimate nodes’s mixed memberships under the proposed model effectively. Our work builds upon and extends the research conducted in [28], [1], and [8], transitioning from non-overlapping multi-layer undirected networks, non-overlapping multi-layer directed networks, and overlapping single-layer undirected networks to overlapping multi-layer directed networks, respectively.
@@ 2.3 Connections to previous work
To demonstrate the effectiveness of our proposed method, we theoretically reveal its consistent estimation properties. Our theoretical framework leverages the Bernstein inequality from [55] and Theorem 4.2 from [56] to derive the row-wise eigenspace error for two debiased aggregation matrices. Subsequently, we employ theoretical frameworks from [8] to establish row-wise error bounds. By doing so, we broaden the scope of the theoretical analysis in [8], extending it from single-layer undirected networks to multi-layer directed networks, and from adjacency matrices to debiased aggregation matrices.
@@ 3 Model framework: multi-layer MM-ScBM
Recall that few previous works have focused on detecting the latent overlapping community structures in multi-layer directed networks, to fill this gap, we need to solve the following fundamental questions:
  • 1.
    Question 1: How can we develop a statistical model to describe/capture/model multi-layer directed networks with latent overlapping community property such that nodes can have mixed community memberships?
  • 2.
    Question 2: Assuming we indeed propose a statistical model to address Question 1, how can we design an efficient algorithm to estimate the true mixed memberships of nodes in multi-layer directed networks generated from this model?
  • 3.
    Question 3: Assuming we indeed devise an algorithm to address Question 2, how can we guarantee that this algorithm enjoys consistent estimation properties under the proposed model? Here, consistent estimation properties indicate that the theoretical upper bound of the error rate of the algorithm decreases to zero as the number of nodes or network layers increases to infinity under this proposed model.
  • 4.
    Question 4: Assuming we theoretically demonstrate the consistent estimation of the proposed algorithm under the proposed model and derive the theoretical upper bound on its error rate, how can we numerically validate the effectiveness of the proposed algorithm?
The rest of this section details the statistical model to address Question 1, while the subsequent three sections present the algorithm, consistency results, and numerical experiments to address Questions 2, 3, and 4, respectively.
@@
This article assumes that the multi-layer directed network $\mathcal{N}$ comprises $L$ distinct layers, each sharing a common set of $n$ interconnected nodes. Let $A_{l}\in\{0,1\}^{n\times n}$ denote the adjacency matrix of the $l$-th layer directed network $\mathcal{N}_{l}$, where $A_{l}(i,j)=1$ indicates the presence of a directed edge from node $i$ to node $j$ (i.e., $i\rightarrow j$) and $A_{l}(i,j)=0$ otherwise, for all $i,j\in[n]$ and $l\in[L]$. Therefore, in $A_{l}$, its $i$-th row captures the outgoing edges from node $i$, while its $i$-th column details the incoming edges to node $i$ at the $l$-th layer for $i\in[n],l\in[L]$. Fig. 1 displays a toy example of a multi-layer directed network, showcasing how nodes establish distinct connections across various layers. For the sending pattern, we assume that all nodes are classified into $K$ distinguishable row communities:
$C_{r,1},C_{r,2},\ldots,C_{r,K}.$ (1)
Let $\Pi_{r}$ be an $n\times K$ row membership matrix common to all layers. Specifically, $\Pi_{r}(i,k)$ represents the “weight” indicating the affiliation of node $i$ to the $k$-th row community $C_{r,k}$ for $i\in[n],k\in[K]$.
@@
$C_{c,1},C_{c,2},\ldots,C_{c,K}.$ (2)
Furthermore, let $\Pi_{c}$ be an $n\times K$ column membership matrix common across all layers and $\Pi_{c}(i,k)$ represents the “weight” indicating the affiliation of node $i$ to the $k$-th column community $C_{c,k}$ for $i\in[n],k\in[K]$.
@@
each community must contain at least one pure node. (3)
Without loss of generality, we rearrange the nodes’ order such that $\Pi_{r}(\mathcal{I}_{r},:)=I_{K\times K}$ and $\Pi_{c}(\mathcal{I}_{c},:)=I_{K\times K}$. Here, $\mathcal{I}_{r}$ represents the indices of $K$ pure nodes, each belonging to a distinct row community, while $\mathcal{I}_{c}$ denotes the indices of $K$ pure nodes, each from a unique column community.
For $l\in[L]$, let $B_{l}$ be a $K\times K$ heterogeneous block probability matrix such that
$B_{l}(k,\tilde{k})\in[0,1]\qquad k\in[K],\tilde{k}\in[K].$ (4)
For each $l$, $B_{l}$ can be asymmetric or symmetric and it captures the community-wise edge probabilities within this layer [28, 1]. In this article, $B_{l}$ can be different for different layers.
For each adjacency matrix $A_{l}$ for $l\in[L]$, the model considered in this article assumes that it is generated from a directed version of MMSB. In particular, we consider the following model.
@@
where $\rho\in(0,1]$.
In the definition of multi-layer MM-ScBM, given that the heterogeneous block probability matrix $B_{l}$ can be different for different layers, this guarantees that $A_{l}$ may be generated from the model with common membership matrices and possibly different edge probabilities. For $i\in[n],j\in[n],l\in[L]$, recall that $\Pi_{r}\in[0,1]^{n\times K},\Pi_{c}\in[0,1]^{n\times K},\|\Pi_{r}(i,:)\|_{1}=1,\|\Pi_{c}(i,:)\|_{1}=1$, and $B_{l}\in[0,1]^{K\times K}$, by basic algebra, we have $0\leq\Pi_{r}(i,:)B_{l}\Pi^{\prime}_{c}(j,:)\leq 1$. Then, under the multi-layer MM-ScBM model defined in Equation (5), we have
$\mathbb{P}(A_{l}(i,j)=1)=\Omega_{l}(i,j)=\rho\Pi_{r}(i,:)B_{l}\Pi^{\prime}_{c}(j,:)\leq\rho,$
$\mathbb{P}(A_{l}(i,j)=0)=1-\Omega_{l}(i,j)=1-\rho\Pi_{r}(i,:)B_{l}\Pi^{\prime}_{c}(j,:)\geq 1-\rho,$
which implies that a reduction in the value of $\rho$ leads to a decrease in the likelihood of an edge forming between any two nodes, thereby controlling the overall sparsity of the multi-layer bipartite network. Consequently, we refer to $\rho$ as the sparsity parameter in this article. Fig. 2 presents a toy example to facilitate a better understanding of Equation (5). In this example, we illustrate the settings of $\Pi_{r}$, $\Pi_{c}$, $\{B_{l}\}^{L}_{l=1}$, and $\rho$. Using Equation (5), we can immediately derive $\{\Omega_{l}\}^{L}_{l=1}$. From this figure, it is straightforward to observe that all elements of $\Omega_{l}$ lie within the range $[0,1]$ for ł$\l\in[L]$, given that $\Omega_{l}$’s elements represent probabilities. Given the set of $L$ adjacency matrices $\{A_{l}\}^{L}_{l=1}$, our goal is to estimate $\Pi_{r}$ and $\Pi_{c}$ in this article. Similar to the consistency analysis conducted in [48, 8, 52, 57] for single-layer networks, we will permit $\rho$ to approach zero as $n$ and $L$ increase in our theoretical analysis.
Refer to captionRefer to captionRefer to caption
Figure 2: A toy example for the settings of $\Pi_{r}$, $\Pi_{c}$, and $\{B_{l}\}^{L}_{l=1}$. The sparsity parameter $\rho$ is set as 1 here. In this example, we consider a 3-layer directed network with 5 nodes belonging to 3 row (and column) communities with different weights. The row and column membership matrices are $\Pi_{r}$ and $\Pi_{c}$, respectively. One can easily check that $\Omega_{l}=\rho\Pi_{r}B_{l}\Pi^{\prime}_{c}$ for $l\in[L]$.
@@
Our model encompasses numerous previous models as specific cases. When all nodes are pure, it simplifies to the multi-layer ScBM discussed in [1]. If the network is undirected and all nodes are pure, it reduces to the multi-layer SBM studied in [10, 18, 17, 28]. Further, for $L=1$, it degenerates to the multi-way blockmodel presented in [49]. In the case of $L=1$ and an undirected network, it reduces to the MMSB model proposed in [5]. Moreover, when all nodes are pure and the network only has a single layer, it reduces to the ScBM model described in [9]. Finally, if the network is undirected, it simplifies to the SBM introduced in [44].